please derive the binary recurrence equation ie
t(n) = t(n/2) + 1,
t(1)=1
given that n is not restricted to be power of two by considering the case that n can either be an odd or even number.
Answer #1

Here we have to solve the given recurrence relation ,

Using substitution method

1. Putting n = n /2 .

2. Finding Relation of T(n), for n = n/2.

3. Generalising T(n) for k steps.

4. After constantly dividing by 2 at last we will get 1.

So , equate

5. As , , so T() = T(1) , which is equal to 1.

6. At last substitute all values in generalised equation of T(n).

Final answer will be , i.e. big O of .

It is valid for all n greater than 1.

Please upvote.